OpenStudy (anonymous):

Calculus. {arctan 2n}

OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit.

OpenStudy (anonymous):

I know that arctan is bounded into (-pi/2 , pi/2). Does this mean that the answer is one of these?

OpenStudy (anonymous):

That means your graph will lie in 1st and 4th quadrant only

OpenStudy (anonymous):

\[f(n)=\tan^-1(2n)\]

OpenStudy (anonymous):

so as n approached infinity, what does arctan do?

OpenStudy (anonymous):

if n approaches infinity, so does 2n but we all know \[\tan(\frac{\pi}{2})=\]

OpenStudy (anonymous):

i know that as n approaches zero, arctan approaches pi/2

OpenStudy (anonymous):

\[\tan(\frac{\pi}{2})=infinity\]

OpenStudy (anonymous):

thus as 2n approaches infinity, arctan approaches pi/2

OpenStudy (anonymous):

as n approaches 0, arctan approaches 0 and not pi/2

OpenStudy (anonymous):

ok i was thinking this: 

OpenStudy (anonymous):

but that is an angle i guess

OpenStudy (jhannybean):

So your question asks if \(a_n = \tan^{-1}(2n)\) converges or diverges?

OpenStudy (anonymous):

yes and if it converges, what is the limit

OpenStudy (anonymous):

nishant told me it is pi/2 and he i correct, cuz i checked... i was just hoping to get a visual of how that works is all

OpenStudy (jhannybean):

So doesnt the graph of inverse tangent just look like